OpenStudy (anonymous):

Given the equation y − 3 = 1/2(x + 6) in point-slope form, identify the equation of the same line in standard form.

OpenStudy (anonymous):

Standard Form is given by: \[\large ax + by + c = 0\] Multiply both the sides by 2, then: \[2y -3 = x + 6\] Or, \[x - 2y = -3-6\] \[x - 2y + 9 = 0\]